OpenStudy (anonymous):

if g(x)=3(x-2), find the value of x if g(x)=6

OpenStudy (zepp):

What you have to do it to plug g(x)'s value into your equation 6 = 3(x-2) Then solve :)

OpenStudy (anonymous):

so 6=3x-6 than you add 6 t0 both sides and you get 12=3x then divide to get 4=x?

OpenStudy (anonymous):

yep, that's ALL!

OpenStudy (zepp):

Yep, that's right, or you can 6 = 3(x-2) 6/3 = x-2 2 = x -2 x = 2 + 2 x = 4 But either way is fine ;D

OpenStudy (anonymous):

Thank you so much !!!!!

OpenStudy (zepp):

Np ;)